OpenStudy (anonymous):

log base 3 4+ log base3 x = 2

OpenStudy (anonymous):

Use this property if log(m) + log(n) = log(m*n) b b b So log(4 * x) = 2 3 Now use log(m) = x then m = b^x b So we will get 4x = 3^2 ==> 4x= 9 ===> x = 9/4